Please help me with this problem. Find the percent of change in price if the old selling price was $20 and the new selling price is $15.
Answer by nerdybill(7384)

Please help me with this problem. Find the percent of change in price if the old selling price was $20 and the new selling price is $15.
.
15/20 * 100 = 0.75 * 100 = 75%
.
So, you could think of it as:
75% of the original price